# Spin-2 fields and cosmic censorship

### v1

@@ Abstract
We evaluate the possibility of the formation of naked singularities in the interaction of Kerr black holes with test massless spin-2 fields. We analyse the scattering problem for both extremal and nearly extremal black holes by incorporating the explicit form of the absorption probabilities. We show that extremal black holes become non-extremal after the interaction with the modes with both positive and negative absorption probabilities. For nearly extremal black holes the absorption probability for the challenging modes turns out to be of the order $\epsilon^{5}$, where $\epsilon$ parametrizes the closeness to extremality. This value implies that the event horizon is preserved without the need to employ the backreaction effects. The result directly follows from the incorporation of the absorption probabilities into the analysis. We also review and extend our previous comments on Sorce-Wald method. We show that the order magnitude problems can be avoided by abandoning the non-physical parameter $\lambda$. The conditions derived by Sorce-Wald can be legitimately used to incorporate the back-reaction effects by avoiding to multiply them by $\lambda^{2}$, thereby preserving the magnitude of their contribution. We apply this method to the problems of over-charging R-N black holes and over-spinning Kerr black holes, in a rough analysis ignoring absorption probabilities which could apply to test bodies.
@@ I Introduction
A generic feature of general relativity is the development of singularities as a result of gravitational collapse pensing . The collapse is expected to end up in a black hole surrounded by an event horizon which is a one way membrane disabling the causal connection between the singularity and the space-time outside. If the distant observers do not encounter the singularities or any effects propagating out of them, the smooth causal structure of the spacetime can be preserved at least outside the black hole region bounded by the event horizon. The weak form of the cosmic censorship conjecture asserts that singularities should always be covered by event horizons which causally disconnects them from the space-time outside ccc . A concrete proof of the cosmic censorship conjecture appears to be elusive.
@@
The smooth structure of spacetimes would also be disrupted if it was possible to destroy event horizons and expose the singularity to outside observers in the interactions of black holes with test particles and fields. In the first of the thought experiments to check this possibility, Wald has shown that particles carrying sufficiently large angular momentum or charge to overspin/overcharge Kerr-Newman black holes are not absorbed by the black holes wald74 . In the following decades many similar thought experiments were constructed which involve test particles and fields hu ; js ; backhu ; backjs ; f1 ; gao ; siahaan ; magne ; yuwen ; higher ; v1 ; he ; wang ; jamil ; shay3 ; shay4 ; zeng ; semiz ; q1 ; q2 ; q3 ; q4 ; q5 ; q6 ; q7 ; overspin ; emccc ; natario ; duztas2 ; mode ; taubnut ; kerrsen ; hong ; yang ; bai ; tjphys ; khoda ; ong . The analysis has also been extended to the asymptotically de-Sitter and anti-de Sitter cases btz ; gwak3 ; chen ; ongyao ; mtz ; he2 ; dilat ; yin ; btz1 ; yang1 ; shay5 ; gwak4 ; shay6 ; corelli ; sia2 . It turns out that cosmic censorship conjecture remains valid for perturbations satisfying the null energy condition, employing the backreaction effects if necessary. We have derived some counter-examples in alternative theories of gravity like Kerr-MOG black holes kerrmog . However, this points to the existence of some caveats in the underlying theory rather than the invalidity of the conjecture.
@@
The validity of the cosmic censorship conjecture and the laws of black hole dynamics for perturbations satisfying the null energy equation relies on the fact that there exists a lower bound for the energy of the perturbation to allow its absorption by the black hole. For test fields this lower bound corresponds to the superradiance limit where the absorption probability vanishes. One cannot derive an analogous limit for fermionic fields the energy momentum tensor of which does not satisfy the null energy condition. The absorption of the low energy modes is allowed which leads to problems regarding the validity of cosmic censorship and the laws of black hole dynamics duztas ; toth ; generic ; spinhalf ; threehalves . The analyses for the cases that do and do not satisfy the null energy condition are fundamentally different and the results do not imply each other. (See threehalves for a detailed comparison of the perturbations that do and do not satisfy the null energy condition.)
@@
In this work, we evaluate the interaction of massless spin-2 fields with Kerr black holes to search for tailored modes of the field that can overspin the black hole into a naked singularity. The energy momentum tensor for spin-2 fields satisfies the null energy condition. There exists a lower limit for the energy/frequency of a test field to allow its absorption by the black hole. The modes with a lower energy/frequency are reflected back to infinity with a larger amplitude; i.e. they exhibit superradiant scattering. Equivalently the absorption probability becomes zero and negative as the frequency decreases. Here, we incorporate the explicit form of the absorption probabilities into the analysis. We use the absorption probabilities derived by Page in his seminal work page . In a rough analysis we would assume that the test field is entirely absorbed if the frequency $\omega$ is larger than the superradiance limit $m\Omega$. However, only a small fraction of a test field with $\omega\gtrsim m\Omega$, is absorbed by the black hole. This fraction approaches zero as the frequency approaches the superradiance limit. In particular for spin (2) fields, as $\omega$ approaches $m\Omega$, the absorption probability decreases as $(\omega-m\Omega)^{5}$. Therefore the incorporation of the absorption probabilities fundamentally changes the course of the analysis for spin (2) fields. There exist modes that could destroy the event horizon if they were entirely absorbed. This destruction would be fixed by the back-reaction effects. However a negligible fraction of the challenging modes is absorbed by the black hole. We show that the event horizon is preserved without invoking the back-reaction effects, when the absorption probabilities are taken into consideration.
@@
Recently Sorce and Wald developed an alternative method and claimed to bring an ultimate solution to the problem of cosmic censorship for perturbations satisfying the null energy condition w2 . In absorp we have scrutinized their derivation and explicitly demonstrated that it involves order of magnitude errors. The order of magnitude errors in $f(\lambda)$ defined by Sorce and Wald, are manifest when one considers the fact that $\delta M$ is inherently a first order quantity for test particles and fields. To avoid a potential misunderstanding, we have not disputed the fact that cosmic censorship remains valid for perturbations satisfying the null energy condition. Our previous works and the current work on spin-2 fields justify this claim. In this work we extend our comments on Sorce-Wald method. We point out that the conditions derived by Sorce and Wald for the first order and second order perturbations are correct. They can be used to derive and incorporate the back-reaction effects which contribute to the interaction to second order. One can simply avoid the order of magnitude problems by abandoning $f(\lambda)$ defined by Sorce-Wald. In $f(\lambda)$, the contribution of the back-reaction effects is multiplied by $\lambda^{2}$, which renders it fourth order. The correct path to follow is to derive the back-reaction effects from the Sorce-Wald condition, and substitute these contributions without multiplying them by $\lambda^{2}$. Here, we use this method to re-evaluate the overcharging problem previously studied by Hubeny hu , and the overspinning problem previously studied by Jacobson-Sotiriou for test bodies js and Düztaş-Semiz for test fields overspin . We adapt our previous methods to show that back-reaction effects derived from the Sorce-Wald condition can be correctly incorporated and fix the overcharging/overspinning problems, in a rough analysis ignoring the absorption probabilities.
@@ II Overspinning problem and the absorption probabilities
In this section we describe how to incorporate the explicit form of the absorption probabilities into the overspinning problem. We start with a Kerr black hole the mass and the angular momentum parameters of which satisfy:
$M^{2}-J\geq 0$ (1)
We envisage that a test field is incident on the black hole from infinity. The test field is partially absorbed by the black hole and partially gets reflected back to infinity. In the test field approximation, the spacetime settles to a new Kerr solution with modified parameters, after a sufficiently long time. If the final parameters of the spacetime violate the main inequality (1) at the end of the interaction, the event horizon ceases to exist. The spacetime parameters represent a naked singularity which is in causal contact with the distant observers.
The energy $(\delta M)$ and the angular momentum $(\delta J)$ of the test field contribute to the mass and angular momentum parameters of the black hole. In a static and axi-symmetric spacetime, the wave-function can be separated in the form
@@
The contributions to the mass and the angular momentum parameters are related by $(\delta J)=(m/\omega)(\delta M)$ In absorp we pointed out that only the fraction of the test field that is absorbed by the black hole contributes to the space-time parameters. The fact that this fraction can be very small for challenging modes, fundamentally alters the course of the analysis. The fraction that will be absorbed by the black hole is given by the absorption probability of the relevant mode. The final parameters of the space-time take the form
@@
$J_{\rm{fin}}=J+\frac{m}{\omega}\Gamma\delta M$ (2)
where $\Gamma$ is the absorption probability of the test field that is incident on the black hole. The absorption probabilities for a classical scattering problem were derived by Page in a seminal work page . For integer spin test fields, the absorption probability is given by:
@@
An explicit form for the absorption probabilities for spin (2) fields is derived by Page in page , by substituting $l=s=2$ in (3) and keeping only the lowest order terms in $\omega$. (The dominant contribution comes from $l=s$ modes) This implies $\omega\ll m\Omega$ which renders this form of the absorption probability inappropriate for our analysis where we require that the test field is absorbed by the black hole; i.e. $\omega\geq m\Omega$. To derive a general form for the absorption probability for spin-2 fields we simply substitute $l=s=2$ in (3)
@@
Using this form of the absorption probability we may proceed to evaluate the validity of cosmic censorship for extremal and nearly extremal black holes.
@@ III Extremal black holes
We can also evaluate the behaviour of the superradiant modes $\omega<m\Omega=(1/M)$. For these modes the absorption probability is negative ($\Gamma<0$). Again we focus on the last two terms of (10) . This time the second term gives a negative contribution, while the third term gives a positive contribution.
@@ IV Nearly extremal black holes and spin-2 fields
In the extremal case the angular velocity of the horizon is $\Omega=(1/2M)$. In this case the superradiance limit ($\omega=m\Omega$) coincides with the upper limit of frequencies $\omega=(1/M)$, that can overspin the black hole. The range of frequencies that can be used to overspin the black hole vanishes. For nearly extremal black holes, there exists a range of frequencies ($m\Omega<\omega<(1/M)$, that could lead to overspinning. The closeness to extremality also determines the closeness of the angular velocity to the critical value $\Omega=(1/2M)$. Therefore the range of frequencies that could lead to overspinning has width $\epsilon$, which also parametrizes the closeness to extremality. The width of the range determines the maximum value of the absorption probability for the modes that would be absorbed by the black hole and lead to the destruction of the event horizon, ignoring the backreaction effects. We showed that the absorption probability for the challenging modes is of the order $\epsilon^{5}$, which implies that $M_{\rm{fin}}^{2}-J_{\rm{fin}}$ is always positive. Had we also ignored the explicit form of the absorption probabilities and only avoided the superradiant range, we would have derived that the test fields with frequency in the range given in (20) could lead to the overspinning of the black hole. For $\Gamma\sim 1$, $\delta M=M\eta\sim M\epsilon$, $M_{\rm{fin}}^{2}-J_{\rm{fin}}$ would vanish to first order for $\omega=(1/M)$, and it would be negative for lower values. However, this overspinning would not be generic. We would be able to fix it by employing the backreaction effects, such as the self energy of the test field and the induced increase in the angular velocity of the black hole which were suggested by Will will . Both of these effects would work in favour of cosmic censorship. However, by incorporating the explicit form of the absorption probabilities, we showed that the event horizon is preserved without employing backreaction effects.
@@ V Sorce-Wald method
In this section we review our previous comments on Sorce-Wald method w2 , which first appeared in absorp . We start with a brief review of subject. Previously, Hubeny hu and Jacobson-Sotiriou js attempted overcharge/overspin nearly extremal black holes by test bodies, in their well-known works. We also evaluated the same problem in a relatively well-known work involving test fields overspin . In these thought experiments, one starts with a nearly extremal Reissner-Nordström black hole parametrized in the form $M^{2}-Q^{2}\sim M^{2}\epsilon^{2}$ for the overcharging process, and a nearly extremal Kerr black hole parametrized as $M^{2}-a^{2}\sim M^{2}\epsilon^{2}$ for the overspinning process. Then one sends in a test particle or field from infinity with respective magnitudes $\delta Q\sim Q\epsilon$ and $\delta M\sim M\epsilon$. In these works the same small parameter was used to parametrize the closeness to extremality and the magnitude of the perturbation. Physically, such a choice is plausible provided that we stay in the test particle/field limit. In these works, the result is that though extremal black holes cannot, nearly extremal black holes can be overcharged or overspun neglecting backreaction effects. The overcharging and overspinning processes observed in these examples are not generic. Quantitatively one finds
@@
which suggests that the negativeness of the right hand sides can be fixed by backreaction effects which contribute to second order in $\delta Q$ and $\delta M$. With the parametrization $\delta Q\sim Q\epsilon$ and $\delta M\sim M\epsilon$, the backreaction effects contribute to second order in $\epsilon$. The backreaction effects involve the self energy of the test particle or field, the induced increase in the angular velocity of the event horizon or different effects depending on the particular problem. Indeed, it was shown that backreaction effects can fix the problem of overcharging and overspinning of nearly extremal black holes by test bodies backhu ; backjs . Recently we have also shown that the induced increase in the angular velocity of the event horizon prevent the overspinning of Kerr black holes by test scalar fields kerrmog .
@@
Actually they have re-derived a well known relation in black hole physics. The first derivation known to this author is by Needham in 1980 needham . The same condition was also derived by Natario, Queimeda, and Vicente in a study involving test fields in 2016 natario . Needham’s condition (25) determines the lower limit for the energy of perturbations satisfying the null energy condition. It identically leads to the same results derived by Hubeny and Jacobson-Sotiriou for the lower bound of the energy in the Reissner-Nordström and Kerr cases, who used different methods to derive the results. For bosonic test fields with energy $\delta M$ and angular momentum $\delta J=(m/\omega)\delta M$, it precisely gives the superradiance condition $\omega\geq m\Omega$.
@@
The claim of Sorce and Wald is that $f(\lambda)$ becomes negative for the first order terms in $\lambda$, then it becomes positive again due to the contribution of the second order terms in $\lambda$, in accord with the previous results. However, a simple order of magnitude analysis reveals that $f(\lambda)$ is simply equal to $M^{2}\epsilon^{2}$ to second order. Contrary to the claim of Sorce and Wald, the terms that are first order in $\lambda$ cannot make $f(\lambda)$ negative since their contribution is of the order $(\lambda\epsilon^{2})$, and the terms that are second order in $\lambda$ cannot fix anything as their contribution is actually of the order $O(\lambda^{2}\epsilon^{2})$. Apparently $f(\lambda)$ defined by Sorce and Wald does not reproduce the previous results by Hubeny, Jacobson-Sotiriou, and Düztaş-Semiz. This is the order of magnitude problem we referred to in absorp . The order of magnitude problem in $f(\lambda)$ is a brute algebraic fact, not subject to controversy.
The main contribution of Sorce and Wald in w2 is the derivation of the Sorce-Wald condition (26) for the second order perturbations. However, the backreaction effects described by this condition are rendered ineffective as they are multiplied by $\lambda^{2}$ in $f(\lambda)$. To elucidate this point, consider the second order perturbation for a Kerr black hole perturbed by a charged particle, derived in w2 .

### v2

@@ Abstract
We evaluate the possibility of the formation of naked singularities in the interaction of Kerr black holes with test massless spin-2 fields. We analyse the scattering problem for both extremal and nearly extremal black holes by incorporating the explicit form of the absorption probabilities. We show that extremal black holes become non-extremal after the interaction with the modes with both positive and negative absorption probabilities. For nearly extremal black holes the absorption probability for the challenging modes turns out to be of the order $\epsilon^{5}$, where $\epsilon$ parametrizes the closeness to extremality. Though the highest absorption probability for $m=2$ modes pertains to spin-2 fields, its drastically small value implies that the event horizon is preserved without the need to employ the backreaction effects. The result can be qualitatively extrapolated to hypothetical higher integer spin fields. We also review and extend our previous comments on Sorce-Wald method. We show that the order magnitude problems can be avoided by abandoning the non-physical parameter $\lambda$. The conditions derived by Sorce-Wald can be legitimately used to incorporate the back-reaction effects by avoiding to multiply them by $\lambda^{2}$, thereby preserving the magnitude of their contribution. We apply this method to the problems of over-charging R-N black holes and over-spinning Kerr black holes, in a rough analysis ignoring absorption probabilities which could apply to test bodies.
@@ I Introduction
A generic feature of general relativity is the development of singularities as a result of gravitational collapse [1]. The collapse is expected to end up in a black hole surrounded by an event horizon which is a one way membrane disabling the causal connection between the singularity and the space-time outside. If the distant observers do not encounter the singularities or any effects propagating out of them, the smooth causal structure of the spacetime can be preserved at least outside the black hole region bounded by the event horizon. The weak form of the cosmic censorship conjecture asserts that singularities should always be covered by event horizons which causally disconnects them from the space-time outside [2]. A concrete proof of the cosmic censorship conjecture appears to be elusive.
@@
The smooth structure of spacetimes would also be disrupted if it was possible to destroy event horizons and expose the singularity to outside observers in the interactions of black holes with test particles and fields. In the first of the thought experiments to check this possibility, Wald has shown that particles carrying sufficiently large angular momentum or charge to overspin/overcharge Kerr-Newman black holes are not absorbed by the black holes [3]. In the following decades many similar thought experiments were constructed which involve test particles and fields [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41]. The analysis has also been extended to the asymptotically de-Sitter and anti-de Sitter cases [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58]. It turns out that cosmic censorship conjecture remains valid for perturbations satisfying the null energy condition, employing the backreaction effects if necessary. We have derived some counter-examples in alternative theories of gravity like Kerr-MOG black holes [59]. However, this points to the existence of some caveats in the underlying theory rather than the invalidity of the conjecture.
@@
The validity of the cosmic censorship conjecture and the laws of black hole dynamics for perturbations satisfying the null energy equation relies on the fact that there exists a lower bound for the energy of the perturbation to allow its absorption by the black hole. For test fields this lower bound corresponds to the superradiance limit where the absorption probability vanishes. One cannot derive an analogous limit for fermionic fields the energy momentum tensor of which does not satisfy the null energy condition. The absorption of the low energy modes is allowed which leads to problems regarding the validity of cosmic censorship and the laws of black hole dynamics [60, 61, 62, 63, 64]. The analyses for the cases that do and do not satisfy the null energy condition are fundamentally different and the results do not imply each other. (See [64] for a detailed comparison of the perturbations that do and do not satisfy the null energy condition.)
@@
In this work, we evaluate the interaction of massless spin-2 fields with Kerr black holes to search for tailored modes of the field that can overspin the black hole into a naked singularity. The energy momentum tensor for spin-2 fields satisfies the null energy condition. There exists a lower limit for the energy/frequency of a test field to allow its absorption by the black hole. The modes with a lower energy/frequency are reflected back to infinity with a larger amplitude; i.e. they exhibit superradiant scattering. Equivalently the absorption probability becomes zero and negative as the frequency decreases. Here, we incorporate the explicit form of the absorption probabilities into the analysis. We use the absorption probabilities derived by Page in his seminal work [65]. In a rough analysis we would assume that the test field is entirely absorbed if the frequency $\omega$ is larger than the superradiance limit $m\Omega$. However, only a small fraction of a test field with $\omega\gtrsim m\Omega$, is absorbed by the black hole. This fraction approaches zero as the frequency approaches the superradiance limit. In particular for spin (2) fields, as $\omega$ approaches $m\Omega$, the absorption probability decreases as $(\omega-m\Omega)^{5}$. Therefore the incorporation of the absorption probabilities fundamentally changes the course of the analysis for spin (2) fields. There exist modes that could destroy the event horizon if they were entirely absorbed. This destruction would be fixed by the back-reaction effects. However a negligible fraction of the challenging modes is absorbed by the black hole. We show that the event horizon is preserved without invoking the back-reaction effects, when the absorption probabilities are taken into consideration.
@@
Recently Sorce and Wald developed an alternative method and claimed to bring an ultimate solution to the problem of cosmic censorship for perturbations satisfying the null energy condition [66]. In [67] we have scrutinized their derivation and explicitly demonstrated that it involves order of magnitude errors. The order of magnitude errors in $f(\lambda)$ defined by Sorce and Wald, are manifest when one considers the fact that $\delta M$ is inherently a first order quantity for test particles and fields. To avoid a potential misunderstanding, we have not disputed the fact that cosmic censorship remains valid for perturbations satisfying the null energy condition. Our previous works and the current work on spin-2 fields justify this claim. In this work we extend our comments on Sorce-Wald method. We point out that the conditions derived by Sorce and Wald for the first order and second order perturbations are correct. They can be used to derive and incorporate the back-reaction effects which contribute to the interaction to second order. One can simply avoid the order of magnitude problems by abandoning $f(\lambda)$ defined by Sorce-Wald. In $f(\lambda)$, the contribution of the back-reaction effects is multiplied by $\lambda^{2}$, which renders it fourth order. The correct path to follow is to derive the back-reaction effects from the Sorce-Wald condition, and substitute these contributions without multiplying them by $\lambda^{2}$. Here, we use this method to re-evaluate the overcharging problem previously studied by Hubeny [4], and the overspinning problem previously studied by Jacobson-Sotiriou for test bodies [5] and Düztaş-Semiz for test fields [29]. We adapt our previous methods to show that back-reaction effects derived from the Sorce-Wald condition can be correctly incorporated and fix the overcharging/overspinning problems, in a rough analysis ignoring the absorption probabilities.
@@ II Overspinning problem and the absorption probabilities
In this section we describe how to incorporate the explicit form of the absorption probabilities into the overspinning problem. We start with a Kerr black hole the mass ($M$)and the angular momentum ($J$) parameters of which satisfy:
$M^{2}-a^{2}\geq 0$ (1)
where $a=(J/M)$. We envisage that a test field is incident on the black hole from infinity. The test field is partially absorbed by the black hole and partially gets reflected back to infinity. In the test field approximation, the spacetime settles to a new Kerr solution with modified parameters, after a sufficiently long time. If the final parameters of the spacetime violate the main inequality (1) at the end of the interaction, the event horizon ceases to exist. The spacetime parameters represent a naked singularity which is in causal contact with the distant observers. Note that the function $(M^{2}-a^{2})$ becomes positive, negative or zero whenever the simpler function $(M^{2}-J)$ is positive, negative or zero. Therefore it is customary to evaluate the simpler function $(M^{2}-J)$.
The energy $(\delta M)$ and the angular momentum $(\delta J)$ of the test field contribute to the mass and angular momentum parameters of the black hole. In a stationary and axi-symmetric spacetime, the wave-function can be separated in the form
@@
The contributions to the mass and the angular momentum parameters are related by
$\delta J=\frac{m}{\omega}\delta M$ (2)
In [67] we pointed out that only the fraction of the test field that is absorbed by the black hole contributes to the space-time parameters. The fact that this fraction can be very small for challenging modes, fundamentally alters the course of the analysis. The fraction that will be absorbed by the black hole is given by the absorption probability of the relevant mode. The final parameters of the space-time take the form
@@
$J_{\rm{fin}}=J+\delta J=J+\frac{m}{\omega}\Gamma\delta M$ (3)
where $\Gamma$ is the absorption probability of the test field that is incident on the black hole, and we have expressed $\delta J$ in terms of $\delta M$ using (2). The absorption probabilities for a classical scattering problem were derived by Page in a seminal work [65]. For integer spin test fields, the absorption probability is given by:
@@
The contribution of a test field to the angular momentum parameter of the space-time is related to its azimuthal number $m$. Each field encompasses modes $l=0,1,2\ldots$, $m=0,1,2\ldots$. The modes with $m=0$ do not contribute to the angular momentum parameter, and the magnitude of the contribution increases as $m$ increases. Naively, one may ask the question what is the essential difference between a spin-0 and a spin-2 field regarding their contribution to angular momentum, if they can both have $m=2$? The difference lies in the absorption probabilities. For a spin-0 field the highest absorption probability pertains to $l=s=0$ modes. For the modes with $m=1$ ($l\geq 1$), the absorption probability acquires a factor $91/(3)^{2}=1/$ 9. (Check the expression in the square brackets in (4).) For $m=2$ this factor becomes $1/(90)^{2}$. Therefore the contribution of these modes is negligible. However, for spin-2 fields the modes with $l=s=2$ have the highest absorption probability. In this case, these modes become our primary concern, and we let $m=2$ to maximize their contribution to the angular momentum parameter. Actually the highest absorption probability for $m=2$ modes pertains to spin-2 fields. In that respect, one may expect that spin-2 fields are more likely to destroy the event horizon compared to spin-0 and spin-1 case due to the absorption of $m=2$ modes. Here, we check if the absorption of $m=2$ modes of the spin-2 fields lead to overspinning by considering the explicit form of the absorption probabilities.
An explicit form for the absorption probabilities for spin (2) fields is derived by Page in [65], by substituting $l=s=2$ in (4) and keeping only the lowest order terms in $\omega$. (The dominant contribution comes from $l=s$ modes) This implies $\omega\ll m\Omega$ which renders this form of the absorption probability inappropriate for our analysis where we require that the test field is absorbed by the black hole; i.e. $\omega\geq m\Omega$. To derive a general form for the absorption probability for spin-2 fields we simply substitute $l=s=2$ in (4)
@@
Using this form of the absorption probability we may proceed to evaluate the validity of cosmic censorship for extremal and nearly extremal black holes. From our previous works we know that $(\omega-m\Omega)$ must be a small number for overspinning to occur, and the surface gravity $\kappa$ is small for a nearly extremal black hole. We mentioned that the highest absorption probability for $m=2$ modes pertains to spin-2 fields. However, (5) implies that the highest probability involves the fifth power of a small number which may render it negligible.
@@ III Extremal black holes
We can also evaluate the behaviour of the superradiant modes $\omega<m\Omega=(1/M)$. It is well-known that super-radiant modes borrow the excess energy from the rotational parameter of the black hole and drive the black hole away from extremality. Apparently these modes do not challenge the validity of cosmic censorship. We will include them in the analysis just for completeness. For these modes the absorption probability is negative ($\Gamma<0$). Again we focus on the last two terms of (11) . This time the second term gives a negative contribution, while the third term gives a positive contribution.
@@ IV Nearly extremal black holes and spin-2 fields
In the extremal case the angular velocity of the horizon is $\Omega=(1/2M)$. In this case the superradiance limit ($\omega=m\Omega$) coincides with the upper limit of frequencies $\omega=(1/M)$, that can overspin the black hole. The range of frequencies that can be used to overspin the black hole vanishes. For nearly extremal black holes, there exists a range of frequencies ($m\Omega<\omega<(1/M)$, that could lead to overspinning. The closeness to extremality also determines the closeness of the angular velocity to the critical value $\Omega=(1/2M)$. Therefore the range of frequencies that could lead to overspinning has width $\epsilon$, which also parametrizes the closeness to extremality. The width of the range determines the maximum value of the absorption probability for the modes that would be absorbed by the black hole and lead to the destruction of the event horizon, ignoring the backreaction effects. We showed that the absorption probability for the challenging modes is of the order $\epsilon^{5}$, which implies that $M_{\rm{fin}}^{2}-J_{\rm{fin}}$ is always positive. Had we also ignored the explicit form of the absorption probabilities and only avoided the superradiant range, we would have derived that the test fields with frequency in the range given in (21) could lead to the overspinning of the black hole. For $\Gamma\sim 1$, $\delta M=M\eta\sim M\epsilon$, $M_{\rm{fin}}^{2}-J_{\rm{fin}}$ would vanish to first order for $\omega=(1/M)$, and it would be negative for lower values. However, this overspinning would not be generic. We would be able to fix it by employing the backreaction effects, such as the self energy of the test field and the induced increase in the angular velocity of the black hole which were suggested by Will [68]. Both of these effects would work in favour of cosmic censorship. However, by incorporating the explicit form of the absorption probabilities, we showed that the event horizon is preserved without employing backreaction effects.
@@ V Sorce-Wald method
In this section we review our previous comments on Sorce-Wald method [66], which first appeared in [67]. We start with a brief review of subject. Previously, Hubeny [4] and Jacobson-Sotiriou [5] attempted overcharge/overspin nearly extremal black holes by test bodies, in their well-known works. We also evaluated the same problem in a relatively well-known work involving test fields [29]. In these thought experiments, one starts with a nearly extremal Reissner-Nordström black hole parametrized in the form $M^{2}-Q^{2}\sim M^{2}\epsilon^{2}$ for the overcharging process, and a nearly extremal Kerr black hole parametrized as $M^{2}-a^{2}\sim M^{2}\epsilon^{2}$ for the overspinning process. Then one sends in a test particle or field from infinity with respective magnitudes $\delta Q\sim Q\epsilon$ and $\delta M\sim M\epsilon$. In these works the same small parameter was used to parametrize the closeness to extremality and the magnitude of the perturbation. Physically, such a choice is plausible provided that we stay in the test particle/field limit. In these works, the result is that though extremal black holes cannot, nearly extremal black holes can be overcharged or overspun neglecting backreaction effects. The overcharging and overspinning processes observed in these examples are not generic. Quantitatively one finds
@@
which suggests that the negativeness of the right hand sides can be fixed by backreaction effects which contribute to second order in $\delta Q$ and $\delta M$. With the parametrization $\delta Q\sim Q\epsilon$ and $\delta M\sim M\epsilon$, the backreaction effects contribute to second order in $\epsilon$. The backreaction effects involve the self energy of the test particle or field, the induced increase in the angular velocity of the event horizon or different effects depending on the particular problem. Indeed, it was shown that backreaction effects can fix the problem of overcharging and overspinning of nearly extremal black holes by test bodies [6, 7]. Recently we have also shown that the induced increase in the angular velocity of the event horizon prevent the overspinning of Kerr black holes by test scalar fields [59].
@@
Actually they have re-derived a well known relation in black hole physics. The first derivation known to this author is by Needham in 1980 [69]. The same condition was also derived by Natario, Queimeda, and Vicente in a study involving test fields in 2016 [31]. Needham’s condition (26) determines the lower limit for the energy of perturbations satisfying the null energy condition. It identically leads to the same results derived by Hubeny and Jacobson-Sotiriou for the lower bound of the energy in the Reissner-Nordström and Kerr cases, who used different methods to derive the results. For bosonic test fields with energy $\delta M$ and angular momentum $\delta J=(m/\omega)\delta M$, it precisely gives the superradiance condition $\omega\geq m\Omega$.
@@
The claim of Sorce and Wald is that $f(\lambda)$ becomes negative for the first order terms in $\lambda$, then it becomes positive again due to the contribution of the second order terms in $\lambda$, in accord with the previous results. However, a simple order of magnitude analysis reveals that $f(\lambda)$ is simply equal to $M^{2}\epsilon^{2}$ to second order. Contrary to the claim of Sorce and Wald, the terms that are first order in $\lambda$ cannot make $f(\lambda)$ negative since their contribution is of the order $(\lambda\epsilon^{2})$, and the terms that are second order in $\lambda$ cannot fix anything as their contribution is actually of the order $O(\lambda^{2}\epsilon^{2})$. Apparently $f(\lambda)$ defined by Sorce and Wald does not reproduce the previous results by Hubeny, Jacobson-Sotiriou, and Düztaş-Semiz. This is the order of magnitude problem we referred to in [67]. The order of magnitude problem in $f(\lambda)$ is a brute algebraic fact, not subject to controversy.
The main contribution of Sorce and Wald in [66] is the derivation of the Sorce-Wald condition (27) for the second order perturbations. However, the backreaction effects described by this condition are rendered ineffective as they are multiplied by $\lambda^{2}$ in $f(\lambda)$. To elucidate this point, consider the second order perturbation for a Kerr black hole perturbed by a charged particle, derived in [66].